Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: BnA configuration partially done (complete on the 13th pending winds which have interfered with the moves); 4-band dipoles in place; correlator hw resources support up to 6x BlBprs (that many sub-bands or use for additional resolution). Stable software system. New OPT/OST (1.09) deployed today. User Committee this week; Claire giving all (4!) talks for the EVLA; particular interest in data distribution/pipelines/cluster use.
  • Key dates:
    • 07 May - 15 Jun: 4-band science observing
    • 13-31 May: BnA configuration
    • May 11-12, 2011 - NRAO Users' Committee Meeting in Green Bank
    • May 18-20, 2011 - Radio Astronomy & the ISM, Durango
    • May 22-26, 2011 - AAS, Boston
    • May 30, 2011 - NRAO Holiday
  • Staff Travel:
    • LS: 9 May - 20 Jun
    • GM: 25 May - 24 Jun
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul
    • Alison Peck (ALMA): 15 May - 17 May
    • Rieks Jager (ALMA): 16 May - 17 May
    • Bob Sault: 16 May - 23 Jun

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA18 out until 25 May.
    • Receivers: EA05 X band repaired; Ka rcvrs complete (spare in progress); this week focus on EA18 (L, S, X Ku (likely delayed a week)).
    • Calibration (Perley): Ongoing investigation into the switched power compression/expansion; testing this week as possible. Bob to support ferreting out the Zeeman testing.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): 3-bit focus; review from 09May meeting; BlB swaps from rack 8 (reduced total number to 60). Work on development version of CM/CBE (supporting subarrays).
  • Computing (Robnett):

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • 0800-1700 Testing
    • 1800-0900(Tues) Science operations and testing
  • Tue:
    • 0800-1700 Testing/software development time
    • 1800-0630: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags).
  • Correlator: [Clark/Carlson] Delay rocking across bands. Implementation pending factor-of-2 reconciliation.
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week

  • 3-bit samplers (Dhawan+): highest priorities
  • Gain compression/expansion investigation
  • Spectral Line (Linz/Ott): Spectral dynamic range test (TSPE0010.3923557). Complete.
  • Spectral Line (Goss/Chomiuk): BP stability/repeatability (TSPE0005.3699117). Complete..
  • Correlator: Rupen: S-band shadow testing; This week as possible.
  • Correlator: Durand/McMullin/Pokorny/Butler; review of CRC error events (may involve testing).
  • Correlator: Momjian/Rupen: High frequency Zeeman test
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Array: *; Individual issue sleuthing

Commissioning Groups

    • Adv: Low band (Owen):
      • Notes: Status report and readiness. Brisk pace of 2 tests/day; 17 antennas working; 1 flagged; 4 dead; 5 have crossed pol'n; work this week to resolve by Friday's observations.Antennas were ready for Friday's observing; fixed date observations took place throughout the weekend. Reports indicate RFI (not seen at LWA): 1) Most data looks very clean. There are no (?) really strong interfering signals but there are some semi persistent (mostly not seen) but sometimes strong, birdies at: 67.250 MHz (ch 37) 70.000 MHz (ch 81) 77.250 MHz (ch 197) 80.000 MHz (ch 241); bandpass issues.
  • Calibrator Models
    • Notes: Status report and Ka/S update: 32 and 37 GHz observations available for 3C138, 3C147, 3C48 (3C286 by the end of the week); currently the information is put on the web (rather than added directly to CASA/AIPS; after A configuration data is added, it will be added directly. S band is not yet available.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:
  • Imaging (Ott/Golap):
    • Notes:

RSRO/ECSO Reports

RSRO B configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

Demo Science Status

AOB

  • Alison will give a quick summary of ALMA status next meeting.

Misc Actions

  • [McMullin]: Contact to Paul regarding delay in phased array.
  • [Wrobel-Claire]: Identify L-band continuum projects (first step in determining how to educate these users).
  • [Harland]: OPT 4-band frequency settings; fixed date mode. Fixed in 1.09.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
-- JosephMcMullin - 2011-05-09
Topic revision: r4 - 2011-05-10, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback